Determine the amount of kinetic energy absorbed by the brakes when a car is travelling at 60km/hr and brakes to 40km/hr?
1
Expert's answer
2020-06-03T12:01:43-0400

The initial speed of the car is "u=60Km/hr=\frac{50}{3}m/s" ,final speed is "v=40Km/hr=\frac{100}{9}m/s" .

Since, mass of the car is not provided, we assume mass be "m=1Kg"

Let work(energy) done(absorbed) by the brake is "W_b" ,

Thus by law of conservation of energy we get,


"\frac{1}{2}mu^2=\frac{1}{2}mv^2 + W_b\\
\implies W_b=\frac{1}{2}mu^2-\frac{1}{2}mv^2"

Thus on plugin the value we get,


"W_b=\frac{1}{2}\bigg(\frac{50}{3}\bigg)^2-\frac{1}{2}\bigg(\frac{100}{9}\bigg)^2\\
\implies W_b=77.16J"